QUIZ #12: CALCULUS 1A (Stankova) Wednesday, April 21, 2004 Section 10:00–11:00 (Voight)

Name:

Please complete the following problem(s) in the space provided. You may not use a calculator. You will have 15 minutes to complete the quiz.

Please include all relevant intermediate calculations and explain your work when appropriate.

Problem 1. Express the limit as a definite integral on the given interval:

lim_{n→∞} ∑_{i=1}^{n} (e^{(x_{i}^{∗})^{2}} + x_{i}^{∗}\sqrt{1 − x_{i}^{∗}}) Δx, [1, 4].

QUIZ #12: CALCULUS 1A (Stankova) Wednesday, April 21, 2004 Section 11:00–12:00 (Voight)

Name:

Please complete the following problem(s) in the space provided. You may not use a calculator. You will have 15 minutes to complete the quiz.

Please include all relevant intermediate calculations and explain your work when appropriate.

Problem 1. The graph of f is shown. Draw a picture and estimate ∫_{−2}^{4} g(x) dx with three subintervals using:

(a) left endpoints, (b) midpoints.